OpenStudy (anonymous):

what is the sum of the measures of the exterior angles of a decagon? please help and explain

OpenStudy (anonymous):

What is a decagon?

OpenStudy (anonymous):

a polygon with 10sides and 10 angles

OpenStudy (anonymous):

Is it a regular decagon?

OpenStudy (anonymous):

should be. I don't know, im not the one posted this

OpenStudy (anonymous):

http://www.mathopenref.com/decagon.html

OpenStudy (anonymous):

Basically, the sum of the angles inside an n-sided polygon is 180*(n-2)

OpenStudy (anonymous):

a regular polygon*

OpenStudy (anonymous):

not interior, exterior is what i need

OpenStudy (anonymous):

so each angle in the decagon is 144 degrees. To find the measure of each exterior angle, we take advantage of the fact that the sum of the interior and the exterior angle is 180 so the measure of each exterior angle is 180 - 144 = 36 degrees multiply that by 10 to get the sum of the measures of the exterior angles of a regular decagon, which is 360

OpenStudy (anonymous):

for a regular polygon with n sides, each interior angle is\[\frac{180(n-2)}{n}=180-\frac{360}{n}\]degrees. each exterior angle is \[180 - (180-\frac{360}{n})=\frac{360}{n}\]multiply by n to get the sum of the measures of the exterior angles of the n-sides polygon\[360\] therefore, for any n-sided regular polygon, the sum of the measures of the exterior angles will always be equal to 360